# IterComp: Iterative Composition-Aware Feedback Learning from Model Gallery for Text-to-Image Generation

### v1

@@ Abstract
Advanced diffusion models like RPG, Stable Diffusion 3 and FLUX have made notable strides in compositional text-to-image generation. However, these methods typically exhibit distinct strengths for compositional generation, with some excelling in handling attribute binding and others in spatial relationships. This disparity highlights the need for an approach that can leverage the complementary strengths of various models to comprehensively improve the composition capability. To this end, we introduce IterComp, a novel framework that aggregates composition-aware model preferences from multiple models and employs an iterative feedback learning approach to enhance compositional generation. Specifically, we curate a gallery of six powerful open-source diffusion models and evaluate their three key compositional metrics: attribute binding, spatial relationships, and non-spatial relationships. Based on these metrics, we develop a composition-aware model preference dataset comprising numerous image-rank pairs to train composition-aware reward models. Then, we propose an iterative feedback learning method to enhance compositionality in a closed-loop manner, enabling the progressive self-refinement of both the base diffusion model and reward models over multiple iterations. Theoretical proof demonstrates the effectiveness and extensive experiments show our significant superiority over previous SOTA methods (e.g., Omost and FLUX), particularly in multi-category object composition and complex semantic alignment. IterComp opens new research avenues in reward feedback learning for diffusion models and compositional generation.
@@ 3.1 Collecting Human Preferences of Compositionality
For attribute binding, we randomly select 500 prompts from each of the following categories: color, shape, and texture in the T2I-CompBench (Huang et al., 2023). Three professional experts ranked the images generated by the six models for each prompt, and their rankings were weighted to determine the final result. The primary criterion is whether the attributes mentioned in the prompt were accurately reflected in the generated images, especially the correct representation and binding of attributes to the corresponding objects.
@@ 4.1 Experimental Setup
We curate a model gallery of six open-source models, each excelling in different aspects of compositional generation: FLUX (BlackForest, 2024), Stable Diffusion 3 (Esser et al., 2024), SDXL (Betker et al., 2023), Stable Diffusion 1.5 (Rombach et al., 2022), RPG (Yang et al., 2024b), and InstanceDiffusion (Wang et al., 2024a). To ensure the base diffusion model thoroughly and comprehensively learns composition-aware model preferences, we progressively expand the model gallery by incorporating new models (e.g., Omost (Omost-Team, 2024), Stable Cascade (Pernias et al., 2023), PixArt- (Chen et al., 2023)) at each iteration. For performance comparison in compositional generation, we select several state-of-the-art methods, including FLUX (BlackForest, 2024), SDXL (Betker et al., 2023), and RPG (Yang et al., 2024b) to compare with our approach. We use GPT-4o (OpenAI, 2024) for the LLM-controlled methods and to infer the layout from the prompt for the layout-controlled methods.
@@ 4.2 Main Results
User Study
We conducted a comprehensive user study to evaluate the effectiveness of IterComp in compositional generation. As illustrated in fig. 5, we randomly selected 16 prompts for each comparison, and invited 23 users from diverse backgrounds to vote on image compositionality, resulting in a total of 1,840 votes. The results show that IterComp received widespread user approval in compositional generation.
@@ 5 Conclusion
Acknowledgement
The author team would like to deliver sincere thanks to Ruihang Chu from Tsinghua University for his significant suggestions for the refinement of this paper.
@@ Appendix A Appendix
In section A.3, we present the quantitative comparison results between IterComp and other diffusion alignment methods.
In section A.4, we provide more visualization results for IterComp and its base diffusion model, SDXL.
@@ A.2.2 Proof of Theorem 1
A.3 Quantitative Comparison with Other Diffusion Alignment Methods.
We compare IterComp with state-of-the-art diffusion alignment methods, Diffusion-DPO (Wallace et al., 2024) and ImageReward (Xu et al., 2024) in terms of image compositionality and realism. We calculate the average results of these models on T2I-CompBench (Huang et al., 2023), and evaluate image realism via CLIP Score and Aesthetic Score. As demonstrated in table 5, IterComp significantly outperforms previous diffusion alignment methods across all three scenarios. IterComp aggregates composition-aware model preferences from multiple models, which are used to train reward models. Guided by these composition-aware reward models, it achieves comprehensive improvements in compositional generation. Its superior performance in image realism is attributed to the effectiveness of iterative feedback learning, where the self-refinement of both the base diffusion model and reward models across multiple iterations drives significant gains in both compositionality and realism.
A.4 More Visualization Results

### v2

@@ Abstract
Advanced diffusion models like Stable Diffusion 3, Omost, and FLUX have made notable strides in compositional text-to-image generation. However, these methods typically exhibit distinct strengths for compositional generation, with some excelling in handling attribute binding and others in spatial relationships. This disparity highlights the need for an approach that can leverage the complementary strengths of various models to comprehensively improve the composition capability. To this end, we introduce IterComp, a novel framework that aggregates composition-aware model preferences from multiple models and employs an iterative feedback learning approach to enhance compositional generation. Specifically, we curate a gallery of six powerful open-source diffusion models and evaluate their three key compositional metrics: attribute binding, spatial relationships, and non-spatial relationships. Based on these metrics, we develop a composition-aware model preference dataset comprising numerous image-rank pairs to train composition-aware reward models. Then, we propose an iterative feedback learning method to enhance compositionality in a closed-loop manner, enabling the progressive self-refinement of both the base diffusion model and reward models over multiple iterations. Detailed theoretical proof demonstrates the effectiveness of this method. Extensive experiments demonstrate our significant superiority over previous methods, particularly in multi-category object composition and complex semantic alignment. IterComp opens new research avenues in reward feedback learning for diffusion models and compositional generation.
@@ 3.1 Collecting Human Preferences of Compositionality
For attribute binding, we randomly select 500 prompts from each of the following categories: color, shape, and texture in the T2I-CompBench (Huang et al., 2023), resulting in a total of 1,500 prompts. Three professional experts ranked the images generated by the six models for each prompt, and their rankings were weighted to determine the final result. The primary criterion is whether the attributes mentioned in the prompt were accurately reflected in the generated images, especially the correct representation and binding of attributes to the corresponding objects.
@@ 4.1 Experimental Setup
We curate a model gallery of six open-source models, each excelling in different aspects of compositional generation: FLUX (BlackForest, 2024), Stable Diffusion 3 (Esser et al., 2024), SDXL (Betker et al., 2023), Stable Diffusion 1.5 (Rombach et al., 2022), RPG (Yang et al., 2024b), and InstanceDiffusion (Wang et al., 2024a). To ensure the base diffusion model thoroughly and comprehensively learns composition-aware model preferences, we progressively expand the model gallery by incorporating new models (e.g., Omost (Omost-Team, 2024), Stable Cascade (Pernias et al., 2023), PixArt- (Chen et al., 2023)) at each iteration. For performance comparison in compositional generation, we select several state-of-the-art methods, including FLUX (BlackForest, 2024), SDXL (Betker et al., 2023), and RPG (Yang et al., 2024b) to compare with our approach. We use GPT-4o (OpenAI, 2024) for the LLM-controlled methods. Additionally, GPT-4o is also employed to infer the layout from the prompt for the layout-controlled methods.
@@ 4.2 Main Results
We compare IterComp with state-of-the-art diffusion alignment methods, Diffusion-DPO (Wallace et al., 2024) and ImageReward (Xu et al., 2024). As demonstrated in table 5, IterComp significantly outperforms previous diffusion alignment methods across all three scenarios. Iterative feedback learning allows models to achieve self-refinement over multiple iterations, resulting in comprehensive improvements in compositionality and realism.
@@ Appendix A Appendix
In section A.3, we conduct an experimental analysis to assess model stability.
In section A.4, we provide the results of user study.
In section A.5, we present the quantitative comparison between IterComp and RPG.
In section A.6, we present the quantitative comparison between IterComp and two layout-based models: InstanceDiffusion and MIGC.
In fig. 12, we provide more visualization results for IterComp.
@@ A.2.2 Proof of Theorem 1
A.3 Analysis on Model Stability
To evaluate the model stability, we selected five methods for comparison: SD1.5 (Rombach et al., 2022), SDXL (Podell et al., 2023), InstanceDiffusion (Wang et al., 2024a), Diffusion-DPO (Wallace et al., 2024), and FLUX (BlackForest, 2024), along with two evaluation metrics: Complex and CLIP-score. Using the same 50 seeds, we calculated the mean and variance of the models’ performance for these metrics. To facilitate visualization, we used the variance of each method as the radius and scaled it uniformly by a common factor () for stability analysis.
Regarding the stability of compositionality, as shown in fig. 8(a), we found that IterComp not only achieved the best overall performance but also demonstrated superior stability. This can be attributed to the iterative feedback learning paradigm enable the model to analyze and refine its output at each optimization step, effectively self-correcting and self-improving. The iterative training approach enables the model to perform feedback training based on its own generated samples rather than solely relying on external data, this enables the model to steadily improve over multiple iterations based on its own foundation. This enables the model to steadily improve over multiple iterations, building on its existing foundation, which significantly enhances its stability.
For the stability of realism or generation quality, as shown in fig. 8(b), our method also exhibited the highest stability. Therefore, the iterative training approach not only improves the model’s performance but also substantially enhances its stability across different dimensions.
A.4 User Study
We conducted a comprehensive user study to evaluate the effectiveness of IterComp in compositional generation. The study involved 41 randomly selected participants from diverse backgrounds. We compared IterComp with five other methods across four aspects: attribute binding, spatial relationships, non-spatial relationships and overall performance. Each comparison involved 25 prompts, culminating in a final survey of 125 prompts and generating 20,500 votes. From the win rate distribution of IterComp shown in the fig. 9, it is evident that IterComp demonstrates significant advantages across all three aspects of compositional generation.
Specifically, compared to the layout-based model InstanceDiffusion (Wang et al., 2024a), IterComp shows an absolute advantage in attribute binding. For text-based models SDXL (Podell et al., 2023) and FLUX (BlackForest, 2024), IterComp leads significantly in spatial relationships. This highlights that the model gallery design effectively collects composition-aware model preferences and enhances performance across different compositional aspects through iterative feedback learning.
A.5 Comparison between IterComp and RPG
Model | Global | Entity | Attribute | Relation | Other | Average |
IterComp | 89.91 | 88.64 | 86.73 | 84.77 | 89.74 | 81.17 |
RPG | 91.01 | 87.39 | 84.53 | 87.92 | 89.84 | 81.28 |
RPG+IterComp | 92.74 | 91.33 | 89.10 | 92.38 | 90.13 | 84.72 |
Model | Single Obj. | Two Obj. | Counting | Colors | Position | Color Attri. | Overall |
IterComp | 0.97 | 0.85 | 0.63 | 0.86 | 0.33 | 0.41 | 0.675 |
RPG | 0.97 | 0.86 | 0.66 | 0.79 | 0.30 | 0.38 | 0.660 |
RPG+IterComp | 0.99 | 0.90 | 0.72 | 0.90 | 0.35 | 0.48 | 0.723 |
We employed two up-to-date benchmarks: DPG-Bench (Hu et al., 2024) and GenEval (Ghosh et al., 2024) for testing to evaluate the capabilities of IterComp and RPG (Yang et al., 2024b) in compositional generation. As demonstrated in table 6 and table 7, IterComp outperforms RPG in metrics like attributes and colors. This is due to our training of a specific reward model for attribute binding, which iteratively enhances IterComp over multiple iterations. Leveraging the strong planning and reasoning capabilities of LLMs, RPG excels in areas such as relations, counting, and positioning. When IterComp is used as the backbone for RPG, the model exhibits remarkable performance across all aspects. This highlights IterComp’s superiority in compositional generation. It’s important to note that IterComp is a simple SDXL-like model that doesn’t require complex computations during inference. As a result, under the same conditions such as prompts and inference steps, IterComp is nearly three times faster than RPG.
A.6 Comparison between IterComp and Layout-based Methods
We provide additional experiments between IterComp, InstanceDiffusion (Wang et al., 2024a), and MIGC (Zhou et al., 2024). As shown in fig. 10, these examples clearly show that while MIGC and InstanceDiffusion can accurately generate objects in the specified positions of the layout, there is a notable gap in generation quality compared to IterComp, such as aesthetics and details. Moreover, the images generated by these two methods often appear visually unrealistic, with significant flaws such as incomplete violins or mismatches between bicycle and its basket. This highlights the clear superiority of our IterComp on compositional generaton.